A company wants to select no more than 2 projects from a set of 4 possible projects. Which of the following constraints ensures that no more than 2 will be selected, assuming that the P variables are binary and represent whether a project is selected (value of 1) or not (value of 0)?
1
Expert's answer
2020-06-10T19:47:37-0400

Let, the four projects be "x_1,x_2,x_3,x_4"

Hence,The number of ways in which company will select no more than 2 projects out of 4 is equivalent to finding the number of non zero integer solution which satisfy the following constraint,thus

"x_1+x_2+x_3+x_4\leq2"

where,

"x_i\in \{0,1\} \: \forall i\in\{1,2,3,4\}"

To find the number of non zero integer solution, we can introduce a dummy variable "2\geq t\geq 0" such that

"x_1+x_2+x_3+x_4+t=2"

As, each "x_i" has two value,thus it has contribution of "1+x" equally and t contributes "(1+x+x^2)" ,hence we have to find the coefficient of "x^2" in the expansion of "(1+x)^4(1+x+x^2)" will give the number of possible solution and which is

"1+{4\choose 1}+{4\choose 2}=11"